OpenStudy (anonymous):

a hiker in africa discovers a skull that contains 43% of its original amount of c-14 Find the age of the skull to the nearest year N=Noe^-kt No=inital amount of c-14 (at time t=0) N= amount of c-14 at time t K=.0001 ??????????????

OpenStudy (anonymous):

N=0.43No so 0.43No=Noe^-0.0001t 0.43*e^0.0001t=1 Ln 0.43+Lne^0.0001t = Ln 1 Ln0.43 +0.0001t = 0 t=-Ln0,43/0.0001 =8439,7